What is the pH in a saturated solution of Ca (OH)2? (solvent is Pure Water with u 0) 2.
Answer #1

The complete step wise calculation of the pH is given as below:-2 Reaction is given by:- Ca(OH) 2 Ca2+ + 2 COHO - consider solubility of [Ca 2+] =s = [oho Solubility product of Ca (OH) a wiou S = M18 x10-2 m or [OH = 1:18x10-2m Pon = log [ong = -log (1118x102) pon- – log (0.0118) = 1.93 POH = 1.93 since, PH + poH